# Holographic image features of an AdS black hole in Einstein-power-Yang-Mills gravity

### v1

@@ Abstract
By utilizing the AdS/CFT correspondence, we investigate the holographic image of an AdS black hole in Einstein-power-Yang-Mills gravity. The AdS boundary hosts a Gaussian oscillation source, which induces a lensed response on the opposite side of the boundary during propagation through bulk spacetime. The optical system assists observers at the north pole to continuously capture holographic images that show an axisymmetric bright ring known as the Einstein ring. As the observation position shifted, the bright ring gradually transformed into a luminous arc and eventually transitioned into a light point. Simultaneously, we examine the impact of variations in relevant physical quantity on the ring, and present the corresponding brightness curve. The results indicate that as the temperature $T$ and nonlinear Yang Mills charge parameter $q$ increase, the ring radius also increases, while an increase in the chemical potential $u$ leads to a decrease. However, the peak brightness curve of the ring invariably decreases as the values of $T$, $u$, and $q$ increase, albeit to varying degrees. Upon comparing the outcomes of geometric optics, it can be observed that the position of the ring in holography images is consistent with that of the photon sphere.
Xiao-Xiong Zeng,a***E-mail: xxzengphysics@163.com Xin-Yun Hu,bE-mail: huxinyun@126.com Ke-Jian HecE-mail: kjhe94@163.com(Corresponding author)
aDepartment of Mechanics, Chongqing Jiaotong University, Chongqing 400074, China
bCollege of Economic and Management, Chongqing Jiaotong University, Chongqing 400074, China
cDepartment of Physics, College of Computer Science and Electronic Information Engineering, Chongqing Technology and Business University, Chongqing, 400067, China
@@ 1 Introduction
The General Relativity (GR) theory is widely recognized as the dominant gravitational theory because it has been extensively validated through numerous observations [1]. The predictions of GR suggest that black holes, immensely compact celestial objects in our universe, have garnered significant attention from physicists. Recently, the Event Horizon Telescope Collaboration (EHT) has revealed unprecedented images of the black hole situated at the core of radio galaxy M87[2, 3, 4, 5, 6, 7] and the Milky Way’s center Sagittarius A [8, 9, 10, 11, 12, 13], providing compelling evidence for the existence of black holes. One of the most remarkable characteristics exhibited by these images is the presence of a dark region within a luminous ring structure, known as the photon ring and shadow of the black hole, respectively. The strong gravitational effect exerted by the black hole causes the deflection of light in its vicinity, resulting in the formation of a shadow that provides valuable insights into the geometry of spacetime. The study of black hole shadows not only uncovers the fundamental characteristics of black holes, but also serves as a robust tool for constraining various gravitational models [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29].
It is well known that the ($D+1$)-dimensional gravity theory in Anti-de Sitter (AdS) can be associated with the $D$-dimensional Conformal field theory (CFT) due to the AdS/CFT correspondence (Anti-de Sitter spacetime/Conformal field theory), which is also known as holography theory [30, 31, 32]. As a pioneering endeavor, Hashimoto et al. constructed holographic images of the Schwarzschild AdS black hole in the bulk using a given response function of Quantum field theory (QFT) on the boundary within the framework of AdS/CFT duality[33, 34]. The results indicate the presence of a distinct circular structure in two-dimensional spherical thermal holographic images, namely, the Einstein ring. Considering that real quantum materials are typically engineered with a finite chemical potential[35], Liu et al. investigated the holographic image of a charged AdS black hole, where the complex scalar field serves as a probing wave in the bulk spacetime. They observed that the radius of ring remained unaffected by changes in the chemical potential, while it exhibited a distinct dependence on the temperature of the boundary system. Then, the study of a black hole that is dual to a superconductor was also conducted in[36], where the electric current of the superconductor under localized time-periodic external electromagnetic field was considered. They also acquired images that depict the discontinuous changes in the size of the photon ring. Moreover, the investigation into holographic imaging of black holes has been expanded to encompass various gravitational frameworks [37, 38, 39, 40, 41, 42, 43, 44, 45].
@@
Motivated by the aforementioned research, it is justified to construct holographic images of a black hole within the framework of EPYM gravity. Fundamentally, the photon sphere exhibits variations contingent upon the specific bulk dual geometry, thereby inducing alterations of corresponding holographic response. Therefore, we are tempted to investigate the behavior of the lensed response for the black hole in EPYM gravity and construct holographic images, so as to study the potential impacts of the relevant physical quantities on the resulting holographic images. As an essential reference, the alteration of the power exponent $q$ will evidently impact the geometric structure of spacetime. Our concern lies in whether this alteration will be applied to the holographic image and how it will change the features of images. On the other hand, in the background of EPYM gravity, there is a modification to the chemical potential of the boundary system. We delve deeper into investigating whether the chemical potential has distinct implications for holographic images when compared to those arising from the charged black hole in [35]. This could potentially serve as an effective means to differentiate between these two gravitational models.
@@ 2 Construction of holographic images
Due to the phenomenon of gravitational lensing, the emitted light from a distant galaxy can undergo distortion and manifest as an arc or multiple separate images when it passes through an intermediate galaxy along the line of sight. In cases where two galaxies are perfectly aligned, the distorted light forms a ring-shaped structure that surrounds the foreground galaxy, commonly known as an Einstein ring. The Einstein ring serves as one of the most compelling manifestations of GR in the cosmos, providing a unique opportunity to investigate colossal galaxies. If the intermediate galaxy were to be substituted with a black hole, the resulting observed image would inherently yield an abundance of valuable insights regarding the characteristics and properties of the black hole. From this perspective, Hashimoto et al. constructed a holographic Einstein ring of the black hole in Schwarzchild-AdS4 spacetime(Sch-AdS4) based on the AdS/CFT correspondence. In [33, 34], the QFT is considered on $R_{t}\times S^{2}$ and an external source is introduced on $S^{2}$, where the external source is mapped to the boundary condition of the bulk field. The probe wave is emitted from the source located on the AdS boundary, diffracted by the black hole as it traverses through the spacetime of the black hole, and ultimately reaches another point on the AdS boundary, resulting in a lensed response. The asymptotic data of the bulk field corresponds to the response function in relation to an external source, and based on this response function, holographic images of an AdS black hole can be constructed. Consequently, they discovered a correspondence between the resultant Einstein ring and the size of the black hole photon sphere determined through geometric optics calculations.
@@ 4 Holographic Einstein ring of black hole in Einstein-power-Yang-Mills gravity
In Fig. 7, we examine the impact of varying the source frequency $\hat{\omega}$ on the characteristics of the holographic images observed from the north pole of the AdS boundary ($\theta_{obs}=0$), while keeping other relevant parameters fixed at $Q_{YM}=0.5,\xi_{+}=0.5$ and $q=7/8$. The results demonstrate a distinct visibility of the striped pattern, revealing a series closed axisymmetric arrangement of concentric ring. Among them, there is consistently an exceptionally luminous ring, known as the holographic Einstein ring, which corresponds to the photon sphere region of the black hole in terms of geometric optics. The width of the ring decreases and becomes sharper as the value of $\hat{\omega}$ increases.
@@ 4.1 Comparison between holographic and optical results
Regarding the shadow image of a black hole, it is established that the luminous ring depicted in the image corresponds to the location of the photon sphere of the black hole. In order to delve deeper into the physical significance of the holographic Einstein ring, we will analyze the correlation between the bright ring in the image and the photon sphere surrounding a black hole from a geometric optics perspective. We can employ the Eikonal approximation as it transforms the Klein-Gordon equation into the form of Hamilton-Jacobi equation, denoted by
@@
The photons in close proximity to the photon sphere will orbit around the black hole one or more times, resulting in cumulative luminosity within this region. The photons that can be located in the vicinity of the photon sphere also applies to Eq. (46), and the corresponding diagram is shown in Fig. 16.
Figure 16: The ingoing angle and outgoing angle of the photon at the photon sphere.
@@ 5 Conclusions and discussions
Gravitational lensing, a prevalent optical phenomenon in the universe, serves as a paramount research tool and pivotal means within astrophysics. It assumes a momentous role in cosmology encompassing dark matter, dark energy, large-scale gravity, and exoplanet detection. As one of the most vivid demonstrations of gravitational lensing, Einstein rings offer an abundance of valuable insights into the intricate fabric of spacetime. From this perspective, we employ wave optics to investigate the holographic image of an AdS black hole within EPYM gravity framework. The AdS boundary is equipped with an oscillating Gaussian source, and the scalar wave emitted from this source propagates through the bulk spacetime, resulting in a corresponding lensed response on the opposite side of the boundary. The diffraction pattern of the total response function ${\langle\mathcal{O}\rangle}_{\mathcal{J}_{\mathcal{O}}}$ is always present during scalar wave propagation through the black hole, as revealed by extracting the local response function at the boundary. It can be found that the amplitude of ${\langle\mathcal{O}\rangle}_{\mathcal{J}_{\mathcal{O}}}$ exhibits a significant increase with the augmentation of the power exponent $q$. However, an opposite trend is observed as the source frequency $\hat{\omega}$, temperature $T$ and chemical potential $u$ increase.
@@
By utilizing a virtual optical system consisting of a convex lens and a spherical screen ( Note that this optical system is in neither the bulk spacetime nor the boundary), we are able to capture the holographic image of an AdS black hole in EPYM gravity. When the source is positioned at the south pole, the image obtained by observers at the north pole ($\theta_{obs}=0$) appears as a bright axisymmetric ring, i.e., the holographic Einstein ring. The brightness curve of the response function is also provided, where the maximum peak on the curve corresponds to the position of the bright ring. The results indicate that the radius of the ring decreases as the chemical potential $u$ of the boundary system increases. However, this decrease is relatively insignificant at low chemical potentials and becomes more pronounced as the chemical potential reaches higher values, which is different from the effect of chemical potential on ring radius in [35]. Due to the chemical potential of the boundary system is modified under the background of EPYM gravity, the impact of changes in chemical potential on the radius of the resulting ring is emphasized. The impact of boundary system temperature $T$ variation on the ring is contrary to that of chemical potential, that is, the radius of the ring increases when the temperature increases, especially at low temperature. The influence of temperature on the ring radius is consistent with the results in[35].
@@
Considering the dependence of spacetime on the the power exponent $q$, we also investigate the effect of the change of $q$ on the Einstein ring at different observation angles ($\theta_{obs}=0,\pi/4,\pi/2$). When the observation angle is $\theta_{obs}=0$, one can find that the radius of the bright ring in the image is observed to increase with an increasing parameter $q$. When the observation angle is $\theta_{obs}=\pi/4$, the ring structure in the image becomes disrupted, leaving a bright arc visible on the left side of the image, while the right side shows a smaller and less intense arc. The bright arc on the left side remains unaffected as the parameter $q$ increases, but the bright arc on the right side gradually diminishes. When the observation angle is $\theta_{obs}=\pi/2$, the bright arc disappears and only one bright spot is located on the left side of the image. With the increase of parameter $q$, the position of the spot tends to move away from the center of the image. Then, the observable brightness of the resulting image will decrease as parameter $q$ increases, regardless of any changes in the observation angles.
@@
We end with comparing the results between the bright ring obtained within the holographic framework and the position of the photon sphere of black hole obtained through geometric optics. The results demonstrate that, across varying temperatures and chemical potentials, the ingoing angles of the photon ring consistently align numerically with those of the Einstein ring. Hence, a correlation can be established between the positions of the photon sphere and the holographic Einstein ring, thereby affirming the reliability of wave optics within the holographic framework. In addition, holographic images can serve as a crucial tool for discerning the geometric characteristics of different black holes, and also providing specific insights into the phenomenological consequences of black hole dynamics, especially when considering fixed wave sources and optical systems.

### v2

@@ Abstract
By utilizing the AdS/CFT correspondence, we investigate the holographic image of an AdS black hole in Einstein-power-Yang-Mills gravity. The AdS boundary hosts a Gaussian oscillation source, which induces a lensed response on the opposite side of the boundary during propagation through bulk spacetime. The optical system assists observers at the north pole to continuously capture holographic images that show an axisymmetric bright ring known as the Einstein ring. As the observation position shifted, the bright ring gradually transformed into a luminous arc and eventually transitioned into a light point. Simultaneously, we examine the impact of variations in relevant physical quantity on the ring, and present the corresponding brightness curve. The results indicate that as the temperature $T$ and nonlinear Yang Mills charge parameter $q$ increase, the ring radius also increases, while an increase in the chemical potential $u$ leads to a decrease. However, the peak brightness curve of the ring invariably decreases as the values of $T$, $u$, and $q$ increase, albeit to varying degrees. Upon comparing the outcomes of geometric optics, it can be observed that the position of the ring in holography images is consistent with that of the photon ring.
Xin-Yun Hu,a***E-mail: huxinyun@126.com Ke-Jian He,bE-mail: kjhe94@163.com Xiao-Xiong Zengc,bE-mail: xxzengphysics@163.com (Corresponding author)
aCollege of Economic and Management, Chongqing Jiaotong University, Chongqing 400074, People’s Republic of China
bDepartment of Mechanics, Chongqing Jiaotong University, Chongqing 400074, People’s Republic of China
cCollege of Physics and Electronic Engineering, Chongqing Normal University, Chongqing 401331, People’s Republic of China
@@ 1 Introduction
[…]
Despite extensive research on the properties of black hole shadows and their associated dynamics, which has provided substantial insights into black hole physics, many mysteries remain to be unraveled using more realistic models. It is well known that the ($D+1$)-dimensional gravity theory in Anti-de Sitter (AdS) can be associated with the $D$-dimensional Conformal field theory (CFT) due to the AdS/CFT correspondence (Anti-de Sitter spacetime/Conformal field theory), which is also known as holography theory [39, 40, 41]. As a pioneering endeavor, Hashimoto et al. constructed holographic images of the Schwarzschild AdS black hole in the bulk using a given response function of Quantum field theory (QFT) on the boundary within the framework of AdS/CFT duality[42, 43]. The results indicate the presence of a distinct circular structure in two-dimensional spherical thermal holographic images, namely, the Einstein ring, and the size of the ring is consistent with the size of the black hole photon ring observed by geometric optics. Considering that real quantum materials are typically engineered with a finite chemical potential[44], Liu et al. investigated the holographic image of a charged AdS black hole, where the complex scalar field serves as a probing wave in the bulk spacetime. They observed that the radius of ring remained unaffected by changes in the chemical potential, while it exhibited a distinct dependence on the temperature of the boundary system. Then, the study of a black hole that is dual to a superconductor was also conducted in[45], where the electric current of the superconductor under localized time-periodic external electromagnetic field was considered. They also acquired images that depict the discontinuous changes in the size of the photon ring. Moreover, the investigation into holographic imaging of black holes has been expanded to encompass various gravitational frameworks [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57]. In these studies, the characteristics of Einstein ring structures across various parameter spaces have been investigated. It has been discovered that, for a given set of wave sources and optical systems, holographic images serve as an effective tool to differentiate between distinct types of black holes.
@@
Motivated by the aforementioned research, it is justified to construct holographic images of a black hole within the framework of EPYM gravity. Fundamentally, the photon ring exhibits variations contingent upon the specific bulk dual geometry, thereby inducing alterations of corresponding holographic response. Therefore, we are tempted to investigate the behavior of the lensed response for the black hole in EPYM gravity and construct holographic images, so as to study the potential impacts of the relevant physical quantities on the resulting holographic images. As an essential reference, the alteration of the power exponent $q$ will evidently impact the geometric structure of spacetime. Our concern lies in whether this alteration will be applied to the holographic image and how it will change the features of images. On the other hand, in the background of EPYM gravity, there is a modification to the chemical potential of the boundary system. We delve deeper into investigating whether the chemical potential has distinct implications for holographic images when compared to those arising from the charged black hole in [44]. This could potentially serve as an effective means to differentiate between these two gravitational models.
@@ 2 Construction of holographic images
Due to the phenomenon of gravitational lensing, the emitted light from a distant galaxy can undergo distortion and manifest as an arc or multiple separate images when it passes through an intermediate galaxy along the line of sight. In cases where two galaxies are perfectly aligned, the distorted light forms a ring-shaped structure that surrounds the foreground galaxy, commonly known as an Einstein ring. The Einstein ring serves as one of the most compelling manifestations of GR in the cosmos, providing a unique opportunity to investigate colossal galaxies. If the intermediate galaxy were to be substituted with a black hole, the resulting observed image would inherently yield an abundance of valuable insights regarding the characteristics and properties of the black hole. From this perspective, Hashimoto et al. constructed a holographic Einstein ring of the black hole in Schwarzchild-AdS4 spacetime(Sch-AdS4) based on the AdS/CFT correspondence. In [42, 43], the QFT is considered on $R_{t}\times S^{2}$ and an external source is introduced on $S^{2}$, where the external source is mapped to the boundary condition of the bulk field. The probe wave is emitted from the source located on the AdS boundary, diffracted by the black hole as it traverses through the spacetime of the black hole, and ultimately reaches another point on the AdS boundary, resulting in a lensed response. The asymptotic data of the bulk field corresponds to the response function in relation to an external source, and based on this response function, holographic images of an AdS black hole can be constructed. Consequently, they discovered a correspondence between the resultant Einstein ring and the size of the black hole photon ring determined through geometric optics calculations.
@@ 4 Holographic Einstein ring of black hole in Einstein-power-Yang-Mills gravity
In Fig.7, we examine the impact of varying the source frequency $\hat{\omega}$ on the characteristics of the holographic images observed from the north pole of the AdS boundary ($\theta_{obs}=0$), while keeping other relevant parameters fixed at $Q_{YM}=0.5,\xi_{+}=0.5$ and $q=7/8$. The results demonstrate a distinct visibility of the striped pattern, revealing a series closed axisymmetric arrangement of concentric ring. Among them, there is consistently an exceptionally luminous ring, known as the holographic Einstein ring, which corresponds to the photon ring region of the black hole in terms of geometric optics. The width of the ring decreases and becomes sharper as the value of $\hat{\omega}$ increases.
@@ 4.1 Comparison between holographic and optical results
Regarding the shadow image of a black hole, it is established that the luminous ring depicted in the image corresponds to the location of the photon ring of the black hole. In order to delve deeper into the physical significance of the holographic Einstein ring, we will analyze the correlation between the bright ring in the image and the photon ring surrounding a black hole from a geometric optics perspective. We can employ the Eikonal approximation as it transforms the Klein-Gordon equation into the form of Hamilton-Jacobi equation, denoted by
@@
The photons in close proximity to the photon ring will orbit around the black hole one or more times, resulting in cumulative luminosity within this region. The photons that can be located in the vicinity of the photon ring also applies to Eq. (46), and the corresponding diagram is shown in Fig. 16.
Figure 16: The ingoing angle and outgoing angle of the photon at the photon ring.
@@ 5 Conclusions and discussions
Gravitational lensing, a prevalent optical phenomenon in the universe, serves as a paramount research tool and pivotal means within astrophysics. It assumes a momentous role in cosmology encompassing dark matter, dark energy, large-scale gravity, and exoplanet detection. As one of the most vivid demonstrations of gravitational lensing, Einstein rings offer an abundance of valuable insights into the intricate fabric of spacetime. From this perspective, we employ wave optics to investigate the holographic image of an AdS black hole within EPYM gravity framework. The AdS boundary is equipped with an oscillating Gaussian source, and the scalar wave emitted from this source propagates through the bulk spacetime, resulting in a corresponding lensed response on the opposite side of the boundary. The diffraction pattern of the total response function ${\langle\mathcal{O}\rangle}_{\mathcal{J}_{\mathcal{O}}}$ is always present during scalar wave propagation through the black hole, as revealed by extracting the local response function at the boundary. The characteristics of the absolute amplitude of the total response function, for various values of the involved parameters, are depicted in Fig.3-Fig.6. These figures demonstrate how wave oscillations change in response to variations in state parameters. From Fig.3, it can be seen that with the increase of the power exponent $q$, the amplitude peak of ${\langle\mathcal{O}\rangle}_{\mathcal{J}_{\mathcal{O}}}$ shows a steady increasing trend. However, as the frequency of the wave source increases, the amplitude peak of ${\langle\mathcal{O}\rangle}_{\mathcal{J}_{\mathcal{O}}}$ diminishes, and the time period of the scalar wave shortens, see Fig.4. Furthermore, the impacts of the finite chemical potential $u$ and horizon temperature $T$ on the amplitude of the total response function are illustrated in Fig.5 and Fig.6. The elevation in temperature $T$ and chemical potential $u$ will progressively diminish the corresponding amplitude of the response function. However, the reduction is more pronounced with variations in temperature $T$.
@@
Although the diffraction pattern of the response function is analyzed, it does not intuitively provide specific information about the black hole. By utilizing a virtual optical system consisting of a convex lens and a spherical screen ( Note that this optical system is in neither the bulk spacetime nor the boundary), we are able to capture the holographic image of an AdS black hole in EPYM gravity. In all cases, irrespective of the specific parameters involved, when the light source is situated at the south pole, an observer located at the north pole ($\theta_{obs}=0$) will invariably perceive an image manifesting as a luminous, axisymmetric ring, commonly referred to as the holographic Einstein ring. Certainly, variations in the relevant parameters exhibit distinct impacts on specific attributes of the ring, such as its size and brightness. In Fig.7, it is observed that as the value of $\hat{\omega}$ increases, the width of the ring becomes progressively narrower and more sharply defined, while the radius of the ring exhibits a decreasing trend. At the same time, the brightness curve of the response function is also provided, where the maximum peak on the curve corresponds to the position of the bright ring, see Fig.8. The results demonstrate that as frequency increases, the position of the maximum peak progressively shifts away from the center of the screen. In other words, the radius of the bright ring expands, which is consistent with the findings presented in Fig.7. Interestingly, in low frequency bands, an increase in frequency leads to an increase in observed brightness, while in high frequency bands, an increase in frequency leads to a decrease in observed brightness. In Fig.11, we analyzed the influence of the chemical potential $u$ on the obtained images of AdS black hole in EPYM gravity. The results indicate that the radius of the ring decreases as the chemical potential $u$ of the boundary system increases. However, this decrease is relatively insignificant at low chemical potentials ($0.50.3<u<$ 0.5) and becomes more pronounced as the chemical potential reaches higher values ($0.90.7<u<$ 0.9), which is different from the effect of chemical potential on ring radius in [44]. Due to the chemical potential of the boundary system is modified under the background of EPYM gravity, the impact of changes in chemical potential on the radius of the resulting ring is emphasized. From Fig.12, it can be observed that the brightness curve of the lensed response on the screen, corresponding to different chemical potentials $u$, shows a gradual decrease in the intensity of the observable ring brightness as $u$ increases. The impact of boundary system temperature $T$ variation on the ring is contrary to that of chemical potential, that is, the radius of the ring increases when the temperature increases, especially at low temperature. The influence of temperature on the ring radius is consistent with the results in[44].
@@
Considering the dependence of spacetime on the the power exponent $q$, we also investigate the effect of the change of $q$ on the Einstein ring at different observation angles ($\theta_{obs}=0,\pi/4,\pi/2$), as shown in Fig.9. As depicted in the leftmost column of Fig.9, when the observation angle is $\theta_{obs}=0$, one can find that the image of the black hole appears as a bright ring with a series of concentric stripes, which correspond to the diffraction behavior of the total response function. Then, the radius of the bright ring in the image is observed to increase with an increasing parameter $q$, and the same conclusion can be obtained from the brightness curve in Fig.10. When the observation angle is $\theta_{obs}=\pi/4$, as illustrated in the middle column of Fig.9, the ring structure in the image becomes disrupted, leaving a bright arc visible on the left side of the image, while the right side shows a smaller and less intense arc. The bright arc on the left side remains unaffected as the parameter $q$ increases, but the bright arc on the right side gradually diminishes. In the rightmost column of Fig.9, the observation angle is $\theta_{obs}=\pi/2$, the bright arc disappears and only one bright spot is located on the left side of the image. With the increase of parameter $q$, the position of the spot tends to move away from the center of the image. Moreover, the observable brightness of the resulting image will decrease as parameter $q$ increases, regardless of any changes in the observation angles.
@@
We end with comparing the results between the bright ring obtained within the holographic framework and the position of the photon ring of black hole obtained through geometric optics. In Fig.18 and 19, The radii of the photon ring (depicted by the blue curve) and the Einstein ring (represented by discrete red dots) are plotted as functions of varying parameter $q$ values. As anticipated, the ingoing angles of the photon ring consistently align numerically with those of the Einstein ring across varying temperatures $T$ and chemical potentials $u$. Hence, a correlation can be established between the positions of the photon ring and the holographic Einstein ring, thereby affirming the reliability of wave optics within the holographic framework. In addition, holographic images can serve as a crucial tool for discerning the geometric characteristics of different black holes, and also providing specific insights into the phenomenological consequences of black hole dynamics, especially when considering fixed wave sources and optical systems. We aspire for these discoveries to furnish a more vivid and intuitive understanding of Einstein rings and their associated phenomena, thereby facilitating future tabletop experiments.